# Frobenius pull-back of parabolic bundles and dormant opers

### v1

@@ 4.3. Parabolic Frobenius pull-back in a global situation
$\mathscr{F}^{\nabla}:=(\mathcal{F}^{\nabla},\vec{\mathpzc{f}}^{\,\nabla},\vec{a}/p^{N})$ (4.23)
@@
Definition 4.4.
We shall refer to $\mathscr{E}^{F}_{\flat}$ as the parabolic Frobenius pull-back of $\mathscr{E}$ (by $F^{(N)}_{X/S}$). Also, we shall refer to $\mathscr{F}^{\nabla}$ as the horizontal parabolic bundle associated to $\mathscr{F}_{\flat}$.
@@
$\mathcal{B}un_{\mathscr{X}^{(N-1)},\vec{a}/p^{N}}\xrightarrow{\sim}\mathcal{D}^{(N)}\text{-}\mathcal{B}un_{\mathscr{X},\vec{a}}^{\psi=0}.$ (4.24)
@@ 4.4. Transitivity of parabolic Frobenius pull-backs
Let $M$ be an integer with $0\leq M\leq N$. In the non-parabolic case, it is immediate that the operation of takings Frobenius pull-backs is transitive, which means $F^{(M)*}_{X/S}F^{(N-M)*}_{X^{(M)}/S}(-)\cong F^{(N)*}_{X/S}(-)$. This section discusses how can we generalize this transitivity to parabolic Frobenius pull-backs.
@@
Now, let $\mathscr{E}:=(\mathcal{E},\vec{\mathfrak{e}},\vec{a}/p^{N})$ be a parabolic bundle classified by $\mathcal{B}un_{\mathscr{X}^{(N)},\vec{a}/p^{N}}$; it associates the parabolic Frobenius pull-back $\mathcal{E}^{F}:=(\mathcal{E}^{F},\nabla^{F},\vec{\mathfrak{e}}^{F},\vec{a})$ by $F_{X/S}^{(N)}$. Also, since $s_{2}^{M}(\vec{a})$ belongs to $\prod_{i=1}^{r}\Xi_{m_{i},N-M}^{\leq}$, one obtains the parabolic Frobenius pull-back of $(\mathcal{E},\vec{\mathpzc{f}},s_{2}^{M}(\vec{a})/p^{N-M})$ by $F_{X^{(M)}/S}^{(N-M)}:X^{(M)}\rightarrow X^{(N)}$, which we denote by $\mathcal{E}^{\prime}:=(\mathcal{E}^{\prime},\nabla^{\prime},\vec{\mathfrak{e}}\,^{\prime},s_{2}^{M}(\vec{a}))$. The resulting triple $(\mathcal{E}^{\prime},\vec{\mathfrak{e}}\,^{\prime},s_{1}^{M}(\vec{a})/p^{M})$ specifies an object in the category $\mathcal{B}un_{\mathscr{X}^{(M)},s_{1}^{M}(\vec{a})/p^{M}}$, so it associates the parabolic Frobenius pull-back by $F_{X/S}^{(M)}$; we denote it by ′′:=(′′,′′$\mathscr{E}^{\prime\prime}:=(\mathcal{E}^{\prime\prime},\nabla^{\prime\prime},\vec{\mathfrak{e}}\,^{\prime\prime},s_{1}^{M}(\vec{a}))$.
@@
(′′$(\mathcal{E}^{\prime\prime},\nabla^{F^{\prime\prime}},\vec{\mathfrak{e}}\,^{\prime\prime},\vec{a})\xrightarrow{\sim}(\mathcal{E}^{F},\nabla^{F},\vec{\mathfrak{e}},\vec{a}).$ (4.26)
@@ 4.5. Parabolic Higgs bundles
    $\mathscr{F}_{\sharp}:=(\mathcal{F},\vec{\mathpzc{f}},\vec{\alpha},\theta)$ (4.27)
    consisting of a parabolic bundle $(\mathcal{F},\vec{\mathpzc{f}},\vec{\alpha})$ on $\mathscr{X}$ (as in the resp’d portion of (2.6)) and a Higgs field $\theta:\mathcal{F}\rightarrow\Omega_{X/S}\otimes\mathcal{F}$ such that $\mathrm{Ker}(\varpi^{[j]}_{\mathscr{F},i})$ (where $\mathscr{F}:=(\mathcal{F},\vec{\mathpzc{f}})$) is closed under $\theta$ for every $i=1,\cdots,r$ and $j=1,\cdots,m_{i}$.
@@
    We say that a non-logarithmic parabolic Higgs bundle $\mathscr{F}_{\sharp}:=(\mathcal{F},\vec{\mathpzc{f}},\vec{\alpha},\theta)$ is $p^{N}$-nilpotent if $\theta^{p-1}=0$.
@@
In what follows, we consider the local construction of $p^{N}$-nilpotent bundles using each object in $\mathcal{H}ig_{\mathscr{U}^{(N)},a_{\circ}/p^{N}}^{\mathrm{non}\text{-}\mathrm{log}}$ in the way similar to the argument in § 4.1. Let us choose a lifting $\widetilde{\mathscr{U}}^{(N)}:=(\widetilde{U}^{(N)}/\widetilde{S},\widetilde{\sigma}_{\circ})$ of $\mathscr{U}^{(N)}$ modulo $p^{2}$, and take an object $\mathscr{E}_{\sharp}:=(\mathcal{E},\mathfrak{e}_{\circ},a_{\circ}/p^{N},\theta)$ of $\mathcal{H}ig_{\mathscr{U}^{(N)},a_{\circ}/p^{N}}^{\mathrm{non}\text{-}\mathrm{log}}$, where $\mathfrak{e}_{\circ}$ and $a_{\circ}/p^{N}$ are given by (4.2) and (4.1), respectively. We write $\mathcal{G}^{[j]}:=\sigma_{\circ*}\mathcal{E}_{\circ}^{[j]}$ ($j=0,\cdots,m_{\circ}$) and $\mathscr{E}:=(\mathcal{E},\mathfrak{e}_{\circ},a_{\circ}/p^{N})$. Note that $\theta$ preserves $\mathrm{Ker}(\varpi_{\mathscr{E}}^{[j]})\left(\subseteq\mathcal{E}\right)$ for each $j$, so it induces an $\mathcal{O}_{U}$-linear morphism $\overline{\theta}^{[j]}:\mathcal{G}^{[j]}\rightarrow\Omega_{U/S}\otimes\mathcal{G}^{[j]}$. By applying the functor $C_{\widetilde{\mathscr{U}}^{(N)}}^{-1}(-)$ to $(\mathcal{E},\theta)$ (resp., $(\mathcal{G}^{[j]},\overline{\theta}^{[j]})$), we obtain a vector bundle $\mathcal{H}$ on $U$ (resp., an $\mathcal{O}_{U}$-module $\mathcal{H}^{[j]}$) equipped with a structure of $\mathcal{D}^{(N-1)}_{\circ}$-module $\nabla_{\mathcal{H}}$ (resp., $\nabla_{\mathcal{H}}^{[j]}$). Since $\widetilde{\sigma}_{\circ}$ defines a marked point on $\widetilde{U}^{(N)}$, the construction of the equivalence between the categories (a) and (b) implies that, for every $i=1,\cdots,r$, $\mathcal{H}^{[j]}$ is supported on $\mathrm{Im}(\sigma_{\circ})$ and obtained as the direct image of a vector bundle on $S$ via $\sigma_{\circ}$. The surjection $\varpi_{\mathscr{E}}^{[j]}:\mathcal{E}\twoheadrightarrow\mathcal{G}^{[j]}$ yields a surjection $\check{\varpi}_{\mathscr{E}}^{[j]}:(\mathcal{H},\nabla_{\mathcal{H}})\twoheadrightarrow(\mathcal{H}^{[j]},\nabla_{\mathcal{H}}^{[j]})$. Consider the following composite surjection of $\mathcal{D}^{(N-1)}_{\circ}$-modules:
@@
  • $\nabla_{\mathcal{H},p^{N}\cdot D_{\circ}}$ (resp., ,$\nabla^{[j]}_{\mathcal{H},p^{N}\cdot D_{\circ}}$) denotes the $\mathcal{D}^{(N-1)}_{\circ}$-module structure on $\mathcal{H}(p^{N}\cdot D_{\circ})$ (resp., $\mathcal{H}^{[j]}(p^{N}\cdot D_{\circ})$) extending $\nabla_{\mathcal{H}}$ (resp., $\nabla^{[j]}_{\mathcal{H}}$);
@@ 4.6. Parabolic generalization of O-V correspondence
It follows from the definition of “$\mathcal{F}$” that there exists an inclusion $(\mathcal{H},\nabla_{\mathcal{H}})\hookrightarrow(\mathcal{F},\nabla_{\mathcal{F}})$ (where “$(\mathcal{H},\nabla_{\mathcal{H}})$” is the $\mathcal{D}_{\circ}^{(N-1)}$-module introduced in the previous subsection), and it is verified to be an isomorphism. Thus, this yields, via $C_{\widetilde{\mathscr{U}}^{(N)}}$, an isomorphism $(\mathcal{E},\theta)\xrightarrow{\sim}(\mathcal{E}_{\star},\theta_{\star})$, which specifies an isomorphism $\mathscr{E}_{\sharp}\xrightarrow{\sim}\mathscr{E}_{\star,\sharp}$. ∎
@@ 5. Dormant $\mathrm{GL}_{n}^{(N)}$-opers
In the rest of the present paper, we assume that the pair of nonnegative integers $(g,r)$ satisfies $02g-2+r>$ 0.
@@ 5.1. Definition of a dormant $\mathrm{GL}_{n}^{(N)}$-oper
:=(,,$\mathscr{F}^{\heartsuit}:=(\mathcal{F},\nabla,\{\mathcal{F}^{j}\}_{j=0}^{n}),$ (5.1)
@@
    We shall say that a $\mathrm{GL}_{n}^{(N)}$-oper :=(,,$\mathscr{F}^{\heartsuit}:=(\mathcal{F},\nabla,\{\mathcal{F}^{j}\}_{j})$ is dormant if $\nabla$ has vanishing $p^{N}$-curvature.
@@
  • :=(,,$\mathscr{F}^{\heartsuit}:=(\mathcal{F},\nabla,\{\mathcal{F}^{j}\}_{j=0}^{n})$ is a dormant $\mathrm{GL}^{(N)}_{n}$-oper on $\mathscr{X}_{S^{\prime}}$ whose exponent at $\sigma_{i}$ coincides with the multiset $[a^{[1]}_{i},\cdots,a^{[n]}_{i}]$ for every $i=1,\cdots,r$;
@@
Suppose that we are given an object $(\mathscr{F}^{\heartsuit},\eta)$ of $\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}$, where :=(,,$\mathscr{F}^{\heartsuit}:=(\mathcal{F},\nabla,\{\mathcal{F}^{j}\}_{j=0}^{n})$. Since (5.3) is an isomorphism, there exists a composite isomorphism
@@ 5.2. Relation with dormant $\mathrm{PGL}_{n}^{(N)}$-opers
This subsection describes the relationship between $\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}$ and the moduli stack of dormant $\mathrm{PGL}_{n}^{(N)}$-opers. The precise definition of a dormant $\mathrm{PGL}_{n}^{(N)}$-oper (of prescribed radii) is omitted here, so we refer the reader to  [Wak4, § 5] for details. Note that we will only use the “$N=1$” case in the subsequent discussion (cf. Theorem 6.7, (i)), and that case is comprehensively investigated in  [Wak3].
@@
the set of equivalence classes of elements in $\Xi_{n,N}^{<}$, where the equivalence relation is defined as follows: $(a^{[1]},\cdots,a^{[n]})$ and $(b^{[1]},\cdots,b^{[n]})$ are equivalent if there exists an element $c\in\mathbb{Z}$ such that the two subsets of $\mathbb{Z}/p^{N}\mathbb{Z}$ defined, via the quotient $\mathbb{Z}\twoheadrightarrow\mathbb{Z}/p^{N}\mathbb{Z}$, by the sets $\{a^{[j]}+c\}_{j=1}^{n}$ and $\{b^{[j]}\}_{j=1}^{n}$ coincide. (Note that $\overline{\Xi}_{n,N}^{<}$ may be regarded as a subset of “$\Xi_{n,N}$” introduced in  [Wak4, § 6.3].) This set has a natural projection
@@
classifying dormant $\mathrm{PGL}_{n}^{(N)}$-opers on $\mathscr{X}$ of radii $\rho(\vec{a})$. (That is to say, $\overline{\mathcal{O}}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\rho(\vec{a})}$ is the fiber product of the projection “$\Pi_{\rho(\vec{a}),g,r}:\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\rho(\vec{a}),g,r}\rightarrow\overline{\mathcal{M}}_{g,r}$” introduced in  [Wak4, § 6.3] and the classifying morhism $S\rightarrow\mathcal{M}_{g,r}\subseteq\overline{\mathcal{M}}_{g,r}$ of $\mathscr{X}$. Here, $\mathcal{M}_{g,r}$ denotes the moduli stack of $r$-pointed smooth proper curves of genus $g$ over $k$, and ¯$\overline{\mathcal{M}}_{g,r}$ denotes its Deligne-Mumford compactification). Each dormant $\mathrm{GL}_{n}^{(N)}$-oper induces a dormant $\mathrm{PGL}_{n}^{(N)}$-oper via projectivization, and this assignment yields a morphism of $S$-stacks
@@
The structure morphism $\mathcal{R}_{n,\mathcal{L}}\rightarrow S$ of $\mathcal{R}_{n,\mathcal{L}}$ is quasi-finite, proper, and étale (by the assumption $n<p$). Also, it follows from (5.8) that the assignment (:=(,,$(\mathscr{F}^{\heartsuit}:=(\mathcal{F},\nabla,\{\mathcal{F}^{j}\}_{j}),\eta)\mapsto\mathcal{F}^{n-1}$ gives an $S$-morphism
@@ 5.3. Harder-Narasimhan polygons
Let $(m_{1},\cdots,m_{r})$ be an $r$-tuple of positive integers and $\vec{a}:=(a_{1},\cdots,a_{r})$ an element of $\prod_{i=1}^{r}\Xi_{m_{i},N}^{\leq}$ such that $s_{1}^{1}(\vec{a})\in\prod_{i=1}^{r}\Xi_{m_{i},1}^{\leq}$ and $s_{2}^{1}(\vec{a})\in\prod_{i=1}^{r}\Xi_{m_{i},N-1}^{\leq}$ (cf. (4.25)). Also, let $n$ be a positive integer and $\mathscr{E}:=(\mathcal{E},\vec{\mathfrak{e}},\vec{a}/p^{N})$ a rank-$n$ semistable parabolic bundle on $\mathscr{X}^{(N)}$ of parabolic weights $\vec{a}/p^{N}$. Denote by $\mathscr{E}^{F}:=(\mathcal{E}^{F},\nabla^{F},\vec{\mathfrak{e}}^{F},\vec{a})$ the parabolic Frobenius pull-back of $\mathscr{E}$ by $F^{(N)}_{X/S}$ and by
@@
where the “$<$” follows from the fact that $\{\mathcal{F}^{j}\}_{j}$ is the Harder-Narasimhan filtration (which implies $\mu(\mathcal{F}^{j})>\mu(\mathcal{F})$). This contradicts the semistability assumption on $\mathscr{E}$, and hence, proves the first assertion.

### v2

@@ 4.3. Parabolic Frobenius pull-back in a global situation
$\mathscr{F}^{\nabla}:=(\mathcal{F}^{\nabla},\vec{\mathfrak{f}}^{\,\nabla},\vec{a}/p^{N})$ (4.23)
@@
math 4.4.
[…]
@@
$\mathcal{B}un_{\mathscr{X}^{(N-1)},\vec{a}/p^{N}}\xrightarrow{\sim}\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{X},\vec{a}}^{\psi=0}.$ (4.24)
@@ 4.4. Transitivity of parabolic Frobenius pull-backs
[…]
@@
Now, let $\mathscr{E}:=(\mathcal{E},\vec{\mathfrak{e}},\vec{a}/p^{N})$ be a parabolic bundle classified by $\mathcal{B}un_{\mathscr{X}^{(N)},\vec{a}/p^{N}}$; it associates the parabolic Frobenius pull-back $\mathcal{E}^{F}:=(\mathcal{E}^{F},\nabla^{F},\vec{\mathfrak{e}}^{F},\vec{a})$ by $F_{X/S}^{(N)}$. Also, since $s_{2}^{M}(\vec{a})$ belongs to $\prod_{i=1}^{r}\Xi_{m_{i},N-M}^{\leq}$, one obtains the parabolic Frobenius pull-back of $(\mathcal{E},\vec{\mathfrak{f}},s_{2}^{M}(\vec{a})/p^{N-M})$ by $F_{X^{(M)}/S}^{(N-M)}:X^{(M)}\rightarrow X^{(N)}$, which we denote by $\mathcal{E}^{\prime}:=(\mathcal{E}^{\prime},\nabla^{\prime},\vec{\mathfrak{e}}\,^{\prime},s_{2}^{M}(\vec{a}))$. The resulting triple $(\mathcal{E}^{\prime},\vec{\mathfrak{e}}\,^{\prime},s_{1}^{M}(\vec{a})/p^{M})$ specifies an object in the category $\mathcal{B}un_{\mathscr{X}^{(M)},s_{1}^{M}(\vec{a})/p^{M}}$, so it associates the parabolic Frobenius pull-back by $F_{X/S}^{(M)}$; we denote it by $\mathscr{E}^{\prime\prime}:=(\mathcal{E}^{\prime\prime},\nabla^{\prime\prime},\vec{\mathfrak{e}}\,^{\prime\prime},s_{1}^{M}(\vec{a}))$.
@@
$(\mathcal{E}^{\prime\prime},\nabla^{F^{\prime\prime}},\vec{\mathfrak{e}}\,^{\prime\prime},\vec{a})\xrightarrow{\sim}(\mathcal{E}^{F},\nabla^{F},\vec{\mathfrak{e}},\vec{a}).$ (4.26)
@@ 4.5. Parabolic Higgs bundles
    $\mathscr{F}_{\sharp}:=(\mathcal{F},\vec{\mathfrak{f}},\vec{\alpha},\theta)$ (4.27)
[…]
@@
    We say that a non-logarithmic parabolic Higgs bundle $\mathscr{F}_{\sharp}:=(\mathcal{F},\vec{\mathfrak{f}},\vec{\alpha},\theta)$ is $p^{N}$-nilpotent if $\theta^{p-1}=0$.
@@
[…]
@@
  • $\nabla_{\mathcal{H},p^{N}\cdot D_{\circ}}$ (resp., $\nabla^{[j]}_{\mathcal{H},p^{N}\cdot D_{\circ}}$) denotes the $\mathcal{D}^{(N-1)}_{\circ}$-module structure on $\mathcal{H}(p^{N}\cdot D_{\circ})$ (resp., $\mathcal{H}^{[j]}(p^{N}\cdot D_{\circ})$) extending $\nabla_{\mathcal{H}}$ (resp., $\nabla^{[j]}_{\mathcal{H}}$);
@@ 4.6. Parabolic generalization of O-V correspondence
[…]
@@ 5. Dormant $\mathrm{GL}_{n}^{(N)}$-opers
In the rest of the present paper, we assume that the pair of nonnegative integers $(g,r)$ satisfies 2g2+r>02g-2+r>0.
@@ 5.1. Definition of a dormant $\mathrm{GL}_{n}^{(N)}$-oper
$\mathscr{F}^{\heartsuit}:=(\mathcal{F},\nabla,\{\mathcal{F}^{j}\}_{j=0}^{n}),$ (5.1)
@@
    We shall say that a $\mathrm{GL}_{n}^{(N)}$-oper $\mathscr{F}^{\heartsuit}:=(\mathcal{F},\nabla,\{\mathcal{F}^{j}\}_{j})$ is dormant if $\nabla$ has vanishing $p^{N}$-curvature.
@@
  • $\mathscr{F}^{\heartsuit}:=(\mathcal{F},\nabla,\{\mathcal{F}^{j}\}_{j=0}^{n})$ is a dormant $\mathrm{GL}^{(N)}_{n}-$oper on $\mathscr{X}_{S^{\prime}}$ whose exponent at $\sigma_{i}$ coincides with the multiset $a^{[1]}_{i},\cdots,a^{[n]}_{i}]$ for every $i=1,\cdots,r$;
@@
Suppose that we are given an object $(\mathscr{F}^{\heartsuit},\eta)$ of $\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}$, where $\mathscr{F}^{\heartsuit}:=(\mathcal{F},\nabla,\{\mathcal{F}^{j}\}_{j=0}^{n})$. Since (5.3) is an isomorphism, there exists a composite isomorphism
@@ 5.2. Relation with dormant $\mathrm{PGL}_{n}^{(N)}$-opers
This subsection describes the relationship between $\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}$ and the moduli stack of dormant $\mathrm{PGL}_{n}^{(N)}$-opers. The precise definition of a dormant $\mathrm{PGL}_{n}^{(N)}$-oper (of prescribed radii) is omitted here, so we refer the reader to  [Wak4, § 5] for details. Note that we will only use the “N=1N=1” case in the subsequent discussion (cf. Theorem 6.7, (i)), and that case is comprehensively investigated in  [Wak3].
@@
the set of equivalence classes of elements in $\Xi_{n,N}^{<}$, where the equivalence relation is defined as follows: $(a^{[1]},\cdots,a^{[n]})$ and $(b^{[1]},\cdots,b^{[n]})$ are equivalent if there exists an element $c\in\mathbb{Z}$ such that the two subsets of $\mathbb{Z}/p^{N}\mathbb{Z}$ defined, via the quotient $\mathbb{Z}\twoheadrightarrow\mathbb{Z}/p^{N}\mathbb{Z}$, by the sets $\{a^{[j]}+c\}_{j=1}^{n}$ and $\{b^{[j]}\}_{j=1}^{n}$ coincide. (Note that $\overline{\Xi}_{n,N}^{<}$ may be regarded as a subset of $\Xi_{n,N}$” introduced in  [Wak4, § 6.3].) This set has a natural projection
@@
classifying dormant $\mathrm{PGL}_{n}^{(N)}$-opers on $\mathscr{X}$ of radii $\rho(\vec{a})$. (That is to say, $\overline{\mathcal{O}}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\rho(\vec{a})}$ is the fiber product of the projection $\Pi_{\rho(\vec{a}),g,r}:\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\rho(\vec{a}),g,r}\rightarrow\overline{\mathcal{M}}_{g,r}$” introduced in  [Wak4, § 6.3] and the classifying morhism $S\rightarrow\mathcal{M}_{g,r}\subseteq\overline{\mathcal{M}}_{g,r}$ of $\mathscr{X}$. Here, $\mathcal{M}_{g,r}$ denotes the moduli stack of $r$-pointed smooth proper curves of genus $g$ over $k$, and $\overline{\mathcal{M}}_{g,r}$ denotes its Deligne-Mumford compactification). Each dormant $\mathrm{GL}_{n}^{(N)}$-oper induces a dormant $\mathrm{PGL}_{n}^{(N)}$-oper via projectivization, and this assignment yields a morphism of $S$-stacks
@@
The structure morphism $\mathcal{R}_{n,\mathcal{L}}\rightarrow S$ of $\mathcal{R}_{n,\mathcal{L}}$ is quasi-finite, proper, and étale (by the assumption $n<p$). Also, it follows from (5.8) that the assignment $(\mathscr{F}^{\heartsuit}:=(\mathcal{F},\nabla,\{\mathcal{F}^{j}\}_{j}),\eta)\mapsto\mathcal{F}^{n-1}$ gives an $S$-morphism
@@ 5.3. Harder-Narasimhan polygons
Let $(m_{1},\cdots,m_{r})$ be an $r$-tuple of positive integers and $\vec{a}:=(a_{1},\cdots,a_{r})$ an element of $\prod_{i=1}^{r}\Xi_{m_{i},N}^{\leq}$ such that $s_{1}^{1}(\vec{a})\in\prod_{i=1}^{r}\Xi_{m_{i},1}^{\leq}$ and $s_{2}^{1}(\vec{a})\in\prod_{i=1}^{r}\Xi_{m_{i},N-1}^{\leq}$ (cf. (4.25)). Also, let $n$ be a positive integer and $\mathscr{E}:=(\mathcal{E},\vec{\mathfrak{e}},\vec{a}/p^{N})$ a rank-$n$ parabolic bundle on $\mathscr{X}^{(N)}$ of parabolic weights $\vec{a}/p^{N}$ whose parabolic Frobenius pull-back by $F^{(N-1)}_{X/S}$ is semistable. Denote by $\mathscr{E}^{F}:=(\mathcal{E}^{F},\nabla^{F},\vec{\mathfrak{e}}^{F},\vec{a})$ the parabolic Frobenius pull-back of $\mathscr{E}$ by $F^{(N)}_{X/S}$ and by
@@
where the “<<” follows from the fact that $\{\mathcal{F}^{j}\}_{j}$ is the Harder-Narasimhan filtration (which implies $\mu(\mathcal{F}^{j})>\mu(\mathcal{F})$). This contradicts the semistability assumption on $\mathscr{E}$, and hence, proves the first assertion.